A particle starts from rest and accelerates as sho
Answer #1

area of (a-t )graph gives change in velocity

a.

speed(at t=10s) =2*10=20 m/s

speed(at t=20s) =2*10 - 3*5= 5 m/s

b. distance travel in 10 sec = 0.5at2 =0.5*2*100=100m

distance travel in (10 to 15 sec) = ut =20*5=100m

distance travel in ( 15 to 20 sec) = ut +0.5at2 =20*5- 0.5*3*25 =62.5 m

so total distance travel in 20 sec =262.5 m